OpenStudy (anonymous):

The triangle below is isosceles (two congruent sides). Find the value for x. one side say x + 3.8 and the other says 16 and then there is nothing at the bottom.

OpenStudy (anonymous):

The two sides are congruent means equal measures. x +3.8 = 16 Solve for x x = 16 - 3.8 x = 12.2

OpenStudy (anonymous):

the two sides are congruent therefore the equation wil be... x+3.8=16 by transposing or combining like terms...transpose 3.8 to the right side of the eqqquation and we will have... x=16-3.8 i therefore coonclude that the measure of the congruent sides of the isosceles triangle is 12.2